one of the rectangular components of force of 40N is 20N . find the angle it makes with this component
 

Grade:12th pass

2 Answers

Aafiya Shaikh
6 years ago
F(x)=Fcos(alpha) Where alpha is angle between force and that rectangular component So, 20=40cos(alpha) 20/40=cos(alpha) 1/2=cos(alpha) (1/2)cos^-1=alpha alpha=60° I.e. it makes angle of 60° with component
Hrishikesh
4 years ago
Given:
           Resultant force(R) = 40 N
           Force in x-axis direction (A) = 20 N
To find:
           Angle  between "R" vector and "A" vector
Solution:
           A=R cos(x)
           20=40 cos(x)
           20/40= cos(x)
          1/2 = cos(x)
            Cos(60)=cos(x)
So,       x=60°